News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 29 October 1999 2.00pm EST

During the last 48 hours Chandra completed a GTO observation of Sirius-B with the HRC-S and HETG, a series of HRC-I calibration measurements of HR 1099, 2REJ1032+532 and Vega, and a GTO observation of M82 with HRC-I. The Observatory completed a pre-belt ACIS CTI measurement and exited the belts at 1.08pm 10/29. Following post-belt ACIS CTI measurements observations will continue of Capella with the HRC-S and LETG.

Real-time commanding was completed to retract the Low Energy Transmission Grating at 4:16pm on 10/28 after the flight software failed to retract the grating as commanded in the daily load. The reason traced to a potentiometer reading used by the software to determine if the grating is inserted or retracted. The A-side potentiometer reading was higher than the limit value for the "inserted" position, causing the software to take no action (as designed). Data from the last series of grating insertions and retractions show erratic values from the potentiometer. The B-side potentiometer shows nominal behavior consistent with the actual hardware insertions and retractions. Work is underway to assess a software patch to mitigate the behavior and to perform a fault analysis.

The next mission schedule was completed and the load reviewed. The new schedule will continue through 11/6 and contains targets including Hydra-A, NGC3489 and the Coma Cluster.





Return to Status Reports